40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat May 23 22:43:10 2015, ericq, Update, General, ifo recovery log 
    Reply  Wed May 27 02:53:57 2015, ericq, Update, General, ifo recovery log 
Message ID: 11326     Entry time: Wed May 27 02:53:57 2015     In reply to: 11322
Author: ericq 
Type: Update 
Category: General 
Subject: ifo recovery log 

Given my suspicion of fault with the X Green BBPD, Koji generously provided me with another one that he had confirmed to be working. 

However, I turns out I was mistaken. With the existing BBPD, I did indeed witness a beat in the RF output, but it is somehow something like 20dBm smaller than it should be. This is why I missed it the other night. Here's a video of a RF output on a scope, wherein the beat is only barely visible because I've set the trigger level very low. I could not make the beatnote any larger through any alignment motions; I had gotten to this point by doing near/far field overlap on the PSL table. 

I'm not sure what could have caused this. Mode mismatch? By eye, the beam spots looked about the same in the near and far fields, and we haven't had to touch the mode matching in quite some time... I've given up on trying to solve this for tonight. 


Just for kicks, I hooked up the fiber PD IR beatnotes as inputs to the ALS DFD. The X beat is too small to even really see above the control room analyzer's noise floor, but the Y beat looked big enough. With the arms locked on IR, the phase tracker output RMS was a few kHz, so not even worth thinking about any further. Not so surprising. 

Finally, I put back / hooked up everything in its nominal state. 

ELOG V3.1.3-